OpenStudy (anonymous):

What is F(3) ? equation coming soon...

OpenStudy (anonymous):

\[F(x)=\frac{ x ^{2}+5 }{ 4x-3 }\]

sam (.sam.):

Just substitute x=3 \[F(x)=\frac{ x ^{2}+5 }{ 4x-3 } \\ \\ F(3)=\frac{ 3 ^{2}+5 }{ 4(3)-3 } \\ \\ F(3)=?\]

OpenStudy (anonymous):

14/9 ?

sam (.sam.):

Correct

OpenStudy (anonymous):

Thanks!!!